OpenStudy (anonymous):

Use the equation below to determine how many moles of silver nitrate (AgNO3) are needed to produce 6.75 moles of copper (II) nitrate (CU(NO3)2) upon reacting with excess copper (Cu).

OpenStudy (aaronq):

For a general reaction: a A + b B -> c C Make a ratio like this: \(\dfrac{n_A}{a}=\dfrac{n_C}{c}\) For your example it'll be : \(\sf \dfrac{moles~of~AgNO_3}{AgNO_3's ~coefficient}=\dfrac{moles~of~Cu(NO_3)_2}{Cu(NO_3)_2's~coefficient}\) note that you need to have the balanced equation because you need to use the coefficients.